OpenStudy (anonymous):

what is the value of the expression 6x-1[7-2(3x-10)]

OpenStudy (anonymous):

Simplifying 6x + -1[7 + -2(3x + -10)] Reorder the terms: 6x + -1[7 + -2(-10 + 3x)] 6x + -1[7 + (-10 * -2 + 3x * -2)] 6x + -1[7 + (20 + -6x)] Combine like terms: 7 + 20 = 27 6x + -1[27 + -6x] 6x + [27 * -1 + -6x * -1] 6x + [-27 + 6x] Reorder the terms: -27 + 6x + 6x Combine like terms: 6x + 6x = 12x -27 + 12x

OpenStudy (anonymous):

forgot to add this on end when x=5

OpenStudy (anonymous):

a.-27 b.-17 c.13 d.33 e.87

OpenStudy (anonymous):

@LanguageEnthusiast

OpenStudy (anonymous):

\[6\times5-1 (7-2 (3\times5-10))\] Use BODMAS. (Bracket, orders, division, multiplication, addition, subtraction.) \[6×5-1 (7-2 (15-10))\] Subtract 10 from 15. \[6×5-1 (7-2×5)\] \[6×5-1 (7+-10)\] \[6×5-1×-3\] \[30-1 (-3)\] \[30+3\] \[30+3 = 33:\] \[33\]

OpenStudy (anonymous):

oh ok ive never learned bomdas only pemdas

OpenStudy (anonymous):

They are exactly the same, only in Britain we use BODMAS. The question is relatively simple just sub in five where x is, then use BODMAS/PEMDAS to solve for your answer.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem, glad to be of help :-).